OpenStudy (anonymous):

what is the difference between a static sine wave,a pulsating sine wave and a travelling sine wave and what are their mathematical representations?

OpenStudy (anonymous):

@ganeshie8 @mathstudent55

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@UnkleRhaukus @ash2326

OpenStudy (anonymous):

please help me..if I understand this my work will be easy

OpenStudy (unklerhaukus):

A traveling wave is a function of position and time, a standing wave is a function of position (but is constant with time).

OpenStudy (aravindg):

A static sine wave is the usual sine wave we come across in problem. A pulsating one has its amplitude varying with time.

OpenStudy (anonymous):

so in a travelling wave each point experiences every amplitude from the positive maximum to negative maximum?

OpenStudy (unklerhaukus):

yeah for a travelling wave over one cycle, every point experiences the max and min

ganeshie8 (ganeshie8):

static sine wave : \[\large y = A \sin (kx)\]

ganeshie8 (ganeshie8):

travelling sine wave with velocity \(v \), \(\implies \text{displacement = vt}\) : \[\large y = A\sin(k(x-vt))\]

ganeshie8 (ganeshie8):

Clearly the wave is moving right as the transformation is shift right ^^

OpenStudy (anonymous):

@ganeshie8 y=Asin(wt) is this for pulsating sine wave?

OpenStudy (anteater):

So then for a pulsating sine wave A would just be some function that varies with t?

OpenStudy (aravindg):

Indeed!

OpenStudy (anteater):

This will probably confuse things, and I apologize, but I am curious! Would a pulsating wave necessarily be a traveling wave?

OpenStudy (unklerhaukus):

I can imaging a standing wave that is pulsating

OpenStudy (anteater):

Ok, thanks! I was thinking that A could be a constant, or a function of t. Could A also be a function of x and t? I mean if you have a sine function that is a function of two variables, couldn't the Amplitude function also be a function of both variables?

OpenStudy (unklerhaukus):

i guess so